40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 12 15:14:09 2014, manasa, Update, General, X end AUX laser fiber setup ETMXtable.png
    Reply  Mon Dec 22 17:06:14 2014, manasa, Update, General, X end AUX laser fiber setup 
       Reply  Tue Dec 23 13:18:37 2014, manasa, Update, General, X end AUX laser fiber setup 
          Reply  Wed Jan 7 19:49:09 2015, manasa, Update, General, X end space issues ETMXspace.png
             Reply  Thu Jan 29 10:14:21 2015, manasa, Update, General, X end space issues photo.pngdrawing.png
                Reply  Wed Feb 4 21:09:43 2015, manasa, Update, General, X endtable work scheduled tomorrow 
                   Reply  Thu Feb 5 15:21:25 2015, manasa, Update, General, X endtable work  
                      Reply  Tue Feb 10 16:01:21 2015, manasa, Update, General, AUX X fiber coupled 72% AUXxTelescope.png
Message ID: 10834     Entry time: Tue Dec 23 13:18:37 2014     In reply to: 10831     Reply to this: 10873
Author: manasa 
Type: Update 
Category: General 
Subject: X end AUX laser fiber setup 

Quote:

 

Since we will not be doing any major locking, I am taking this chance to move things on the X end table and install the fiber coupler.

The first steering mirror shown in the earlier elog will be a Y1 (HR mirror) and the second one will be a beam sampler (similar to the one installed at the Y endtable for the fiber setup). 

Configuration:

Doubler --> Y1 ---> Lens (f=12.5cm) ---> Beam sampler --->Fiber coupler

The fiber coupler mount will be installed in the green region to the right of the TRX camera. 

This work will involve moving around the TRX camera and the optic that brings the trans image on it.

Let me know if this work should not be done tomorrow morning for any reason.

I was working around the X endtable and PSL table today.

1. Y1 mirror, beam sampler and the fiber coupler have been installed.

2. Removed TRX camera temporarily. The camera will be put back on the table once we have the filter for 532nm that can go with it.

3. Removed an old fiber mount that was not being used from the table. 

4. Lowered the current for X end NPRO while working and put it back up at 2A before closing.

5. The fibers running from the X end to the PSL table are connected at an FC/APC connector on the PSL table. 

6. Found the HEPA left on high (probably from yesterday's work around the PSL table). I have brought it back down and left it that way.

I have not installed the coupling lens as yet owing to the space restrictions - not enough space for footprint of the lens. I have to revisit the telescope design again.

ELOG V3.1.3-